OpenStudy (anonymous):

True or False: (-101 + x) + (101 - x) = 0

OpenStudy (anonymous):

(-101 + x) + (101 - x) = -101 + 101 +x -x = ... ?

OpenStudy (anonymous):

True

OpenStudy (anonymous):

^^ leave a room for imagination!! :P